OpenStudy (anonymous):

use a sum or difference formula to find the exact value of the trigonometric function. cos 165°

OpenStudy (anonymous):

cos165 = cos(180 - 15) = -cos15 is the answer... Remember, cosx where x in 2nd Quadrant between 90 to 180 is negative.. And cos(180-x) = -cosx